Please help! I accidently reset my sons work page and all his answers went away.
Solve the following and show all work
x - 2y = 14
x + 3y = 9


Answer by stanbon(75887)

Solve the following and show all work
x - 2y = 14
x + 3y = 9
----
Subtract the top equation from the bottom and solve for "y":
5y = -5
-------
Divide both sides by 5:
y = -1
------
Solve for "x":
x - 2y = 14
x -2*-1 = 14
x = 12
-----------------
Cheers,
Stan H.
=================